OpenStudy (anonymous):

For each pair of functions f and g below, find f(g(x)) and g(f(x)). Then, determine whether f and g are inverses of each other. Simplify your answers as much as possible. (Assume that your expressions are defined for all x in the domain of the composition. You do not have to indicate the domain.) (a) f(x)=-4x g(x)=1/4x, x neq 0 f(g(x)) g(f(x)) b) f(x)=1/3x, x neq 0 g(x)= 1/3x, x neq 0 f(g(x)) g(f(x))

OpenStudy (gorv):

f(g(x))=f(1/4x)=-4*1/4x=-1/x g(f(x))=g(-4x)=1/4*(-4x)=-1/16x

OpenStudy (gorv):

B f(g(x))=1/(3*1/3x)=x g(f(x))=1/(3*1/3x)=x